OpenStudy (anonymous):

how do you add -pi/4+8pi/4

OpenStudy (aum):

The denominator is the same and so the numerators can be added.\[-\frac{\pi}{4} + \frac{8\pi}{4} = \frac{-\pi + 8\pi}{4} = \frac{7\pi}{4} \]

OpenStudy (anonymous):

thanks a lot!!

OpenStudy (aum):

You are welcome.